# Conditional Wasserstein Distances with Applications in Bayesian OT Flow Matching

### v1

@@ Appendix E Proofs of Section 6
$W_{p,Y}(\mu_{0},\mu_{1})^{p}$ $=\int_{(A\times B)^{2}}\|(y_{1},x_{1})-(y_{2},x_{2})\|^{p}\,\mathrm{d}\alpha=\int_{(A\times B)^{2}}d_{\beta}^{p}((y_{1},x_{1}),(y_{2},x_{2}))\,\mathrm{d}\alpha$
@@
$\geq\int_{B^{2}}\|x_{1}-x_{2}\|^{p}\,\mathrm{d}\pi^{2,4}_{\sharp}\alpha+\beta\int_{A^{2}}\|y_{1}-y_{2}\|^{p}\,\mathrm{d}\pi^{1,3}_{\sharp}\alpha$
@@
In order to proof Proposition 12 we need the following lemma which is a variant of [6, Proposition 7.1.3].
@@
Let $\beta>0$, let $A\subset\mathbb{R}^{d},B\subset\mathbb{R}^{m}$ be compact sets and let $\mu_{n}\to\mu$, $\nu_{n}\to\nu$ in $(\mathcal{P}_{2}(A\times B),W_{2})$. Then there exists a subsequence of optimal plans $\alpha_{n_{k}}$ for $W_{2,d_{\beta}}(\mu_{n_{k}},\nu_{n_{k}})$ and an optimal plan $\alpha\in P_{2}\left((A\times B)^{2}\right)$ for $W_{2,d_{\beta}}(\mu,\nu)$ such that $\alpha_{n_{k}}\to\alpha$ with respect to $\left((\mathcal{P}_{2}\left((A\times B))^{2}\right),W_{2}\right)$.
@@
Let $f:A\times B\to\sqrt{\beta}A\times B$ be defined by $(y,x)\mapsto(\sqrt{\beta}y,x)$. Then for $\mu_{1},\mu_{2}\in P_{2}(A\times B)$ we have that $W_{2,d_{\beta}}(\mu_{1},\mu_{2})=W_{2}(f_{\sharp}\mu_{1},f_{\sharp}\mu_{2})$ since there is a bijection of couplings
@@
$\int_{\left(A\times\sqrt{(}\beta)B\right)^{2}}\|(y_{2},x_{2})-(y_{1},x_{2})\|^{2}\mathrm{d}(f,f)_{\sharp}\alpha=\int_{(A\times B)^{2}}\beta\|y_{2}-y_{1}\|^{2}+\|x_{2}-x_{1}\|\mathrm{d}\alpha$ (94)
which implies that optimal couplings are mapped to optimal couplings. Furthermore
$W_{2}^{2}(f_{\sharp}\mu_{1},f_{\sharp}\mu_{2})$ $=W^{2}_{2,d_{\beta}}(\mu_{1},\mu_{2})\leq\beta W_{2}^{2}(\mu_{1},\mu_{2})$ (95)
$W_{2}^{2}(\mu_{1},\mu_{2})$ $\leq W^{2}_{2,d_{\beta}}(\mu_{1},\mu_{2})=W_{2}^{2}(f_{\sharp}\mu_{1},f_{\sharp}\mu_{2})$ (96)
@@
and thus also $f_{\sharp}\mu_{n}\to f_{\sharp}\mu,f_{\sharp}\nu_{n}\to f_{\sharp}\nu$ in $W_{2}$. Then we can use [6, Proposition 7.1.3] to guarantee the existence of a subsequence of optimal plans $\tilde{\alpha}_{n_{k}}$ for $W_{2}(f_{\sharp}\mu_{n_{k}},f_{\sharp}\nu_{n_{k}})$ such that $\tilde{\alpha}_{n_{k}}\to\tilde{\alpha}$ for an optimal plan $\tilde{\alpha}$ for $W_{2}(f_{\sharp}\mu,f_{\sharp}\nu)$. Thus by (93) there exists a subsequence of optimal plans $\alpha_{n_{k}}$ for $W_{2,d_{\beta}}(\mu_{n},\nu_{n})$ such that $\alpha_{n_{k}}\to\alpha$ for an optimal plan $\alpha$ for $W_{2,d_{\beta}}(\mu,\nu)$. Note that the $W_{2}$ convergence of $\alpha_{n_{k}}\to\alpha$ follows from a computation similar to (95) for $(f,f)_{\sharp}$ instead of $f_{\sharp}$. ∎
Proof of Proposition 12. Since we are in the compact setting the concept of weak and $W_{2}$ convergence coincide which we will use without mentioning in the following [51]. Now by [6, Proposition 7.1.3] and Lemma 17, there exists a subsequence of $\alpha_{n}$ converging in $W_{2}$ to an optimal plan $\alpha^{\beta_{k}}$ for $W_{2,d_{\beta_{k}}}(\mu,\nu)$. Choose $n_{k}$ monotonely increasing such that $W_{2}(\alpha^{\beta_{k}},\alpha_{n_{k}})<\frac{1}{k}$. We know by [29, Proposition 3.11] that $\alpha_{\beta_{k}}\to\alpha$ in $W_{2}$ for an optimal plan $\alpha\in\Gamma^{4}_{Y}(\mu,\nu)$ for $W_{2,Y}(\mu,\nu)$. Thus, for $\epsilon>0$, there exists a $k$ such that $\frac{1}{k}+W_{2}(\alpha^{\beta_{k}},\alpha)<\epsilon$ and we obtain
$W_{2}(\alpha_{n_{k}},\alpha)\leq W_{2}(\alpha_{n_{k}},\alpha^{\beta_{k}})+W_{2}(\alpha^{\beta_{k}},\alpha)\leq\frac{1}{k}+W_{2}(\alpha^{\beta_{k}},\alpha)<\epsilon$
@@
Appendix F Benamou-Brenier like formula for $W_{p,Y}$
Theorem 18.
Let $\mu_{1},\mu_{2}\in\mathcal{P}_{2,Y}(\mathbb{R}^{d}\times\mathbb{R}^{m})$ and denote by $\mathcal{A}(\mu_{1},\mu_{2})$ the set of tuples $(\mu_{t},v_{t})$, where
  • i)
    $\mu_{t}\in\mathcal{P}_{p,Y}$ is a continuous curve,
  • ii)
    There exists a family of disintegrations $\mu_{t}=(\mu_{t})_{y}\otimes P_{Y}$ such that $(\mu_{t})_{y}$ is continuous $P_{Y}$-a.e.,
  • iii)
    $v_{t}$ is a Borel vector field fulfilling $\int_{0}^{1}\|v_{t}\|_{L^{2}(\mu_{t})}\mathrm{d}t<\infty$ and $(v_{t})_{j}=0$ for all $j\leq d$ for $\mu_{t}$ a.e. $(y,x)\in\mathbb{R}^{d+m}$,
  • iv)
    $\mu_{t}$ fulfills the continuity equation in the sense of distributions for $v_{t}$.
Then we have that
$W_{p,Y}(\mu_{1},\mu_{2})=\min_{(\mu_{t},v_{t})\in\mathcal{A}(\mu_{1},\mu_{2})}\left\{\int_{0}^{1}\|v_{t}\|^{2}_{L^{2}(\mu_{t})}\,\mathrm{d}t\right\}$ (97)
In order to prove Theorem 18, we need the following auxiliary lemma.
Lemma 19.
Let $\mu_{t}\in P_{p,Y}(\mathbb{R}^{d}\times\mathbb{R}^{m})$ be a solution of the continuity equation with $v_{t}$ such that $(v_{t})_{j}=0$ for $j\leq d$ and $\int_{0}^{1}\int_{\mathbb{R}^{d+m}}\|v_{t}\|\mathrm{d}\mu_{t}\mathrm{d}t<\infty$. Then the disintegration $(\mu_{t})_{y}$ satisfies the continuity equation for $v_{t,y}$ for $P_{Y}$-a.e. $y\in\mathbb{R}^{d}$.
Proof.
Let $\phi\in C^{\infty}_{c}((0,1)\times\mathbb{R}^{m})$ be a test function and let $g\in C^{\infty}_{c}(\mathbb{R}^{d})$. Then for $\psi(t,y,x)=g(y)\phi(t,x)$ we have that $\psi\in C^{\infty}_{c}((0,1)\times\mathbb{R}^{d+m})$ is a valid test function which we can insert into the continuity equation, use the theorem of Fubini to change the order of the integrals and obtain
$\int_{\mathbb{R}^{d}}g(y)\int_{0}^{1}\int_{\mathbb{R}^{m}}\partial_{t}\phi(x,t)+\langle v_{t,y},\nabla_{x}\phi\rangle\mathrm{d}(\mu_{t})_{y}(x)\mathrm{d}P_{Y}(y)=0$ (98)
Since $g$ was arbitrary we obtain that $\int_{0}^{1}\int_{\mathbb{R}^{m}}\partial_{t}\phi(x,t)+\langle v_{t,y},\nabla_{x}\phi\rangle\mathrm{d}(\mu_{t})_{y}(x)=0$ for $P_{Y}$ a.e. $y\in\mathbb{R}^{d}$. Since $C_{c}^{\infty}((0,1)\times\mathbb{R}^{d})$ contains a dense countable subset in the $\|\cdot\|_{\infty}$ topology we only need to test on countably $\phi$ we can conclude that $(\mu_{t})_{y}$ is a solution of the continuity equation for $v_{t,y}$ for $P_{Y}$ a.e. $y\in\mathbb{R}^{d}$. ∎
Proof of Theorem 18. First we show """"$\leq$"" ≤ ". By Proposition 19, we have that $(\mu_{t})_{y}$ fulfills the continuity equation for $v_{t,x_{1}}$ $P_{Y}$-a.e. and by assumption $(\mu_{t})_{y}$ is weakly continuous. Thus we have that $W_{2}((\mu_{1})_{y},(\mu_{2})_{y})\leq\int_{0}^{1}\int_{\mathbb{R}^{m}}\|v_{t,y}\|^{2}(\mu_{t})_{y}\mathrm{d}t$. Now we get
$W_{2,Y}(\mu_{1},\mu_{2})$ $=\int_{\mathbb{R}^{d}}W_{p}((\mu_{1})_{y},(\mu_{2})_{y})\,\mathrm{d}P_{Y}\leq\int_{\mathbb{R}^{n}}\int_{0}^{1}\int_{\mathbb{R}^{m}}\|v_{t,y}\|^{2}\mathrm{d}(\mu_{t})_{y}\,\mathrm{d}t\mathrm{d}P_{Y}$ (99)
$=\int_{0}^{1}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{m}}\|v_{t,y}\|^{2}\mathrm{d}(\mu_{t})_{y}\,\mathrm{d}P_{Y}\mathrm{d}t=\int_{0}^{1}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{m}}\|v_{t}\|^{2}\mathrm{d}(\mu_{t})_{y}\,\mathrm{d}P_{Y}\mathrm{d}t$ (100)
$=\int_{0}^{1}\|v_{t}\|^{2}_{L^{2}(\mu_{t})}\,\mathrm{d}t,$ (101)
where we used that $\int\|v_{t,y}\|^{2}\mathrm{d}\mu_{t}=\int\|v_{t}\|^{2}\mathrm{d}\mu_{t}$, since $(v_{t})_{j}=0$ for all $j\leq n$ for $\mu_{t}$ a.e. $(y,x)\in\mathbb{R}^{d+m}$.
The direction """"$\geq$"" ≥ " follows from Proposition 6 and Lemma 5ii). $\Box$
Appendix G Implementation Details
We use a setup similar to [50], using the time dependent U-Net architecture from [42] which are trained using Adam [32]. As in [50] we clip the gradient norm to 1 and use exponential moving averaging with a decay of $0.9999$. The differences are we use a constant learning rate of 2e-4, 256 model channels and no dropout. We train using 50k target samples for 300 epochs using a batch size of 500 for the minibatch OT couplings and a batch size of 100 for training the networks. We set the same random seed during training to be able to compare runs for different sources of couplings. The conditional coupling plans are calculated using the Python Optimal Transport package [20]. For inference simulate the corresponding ODEs using the torchiffeq [15] package. To evaluate our results, we use the Fréchet inception distance (FID) [28] [^1] We use the implementation from https://github.com/mseitzer/pytorch-fid.. We compute the distance on 50k training samples, for which we generate 50k samples given the same labels as the training samples.

### v2

@@ Appendix E Proofs of Section 6
$W_{p,Y}(\mu_{0},\mu_{1})^{p}$ $=\int_{(A\times B)^{2}}\|(y_{1},x_{1})-(y_{2},x_{2})\|^{p}\,\mathrm{d}\alpha_{opt}=\int_{(A\times B)^{2}}d_{\beta}^{p}((y_{1},x_{1}),(y_{2},x_{2}))\,\mathrm{d}\alpha_{opt}$
@@
$=\geq\int_{B^{2}}\|x_{1}-x_{2}\|^{p}\,\mathrm{d}\pi^{2,4}_{\sharp}\alpha+\beta\int_{A^{2}}\|y_{1}-y_{2}\|^{p}\,\mathrm{d}\pi^{1,3}_{\sharp}\alpha$
@@
In order to proof Proposition 12 we need some auxiliary results, in particular the following lemma which is a variant of [6, Proposition 7.1.3].
@@
Let $\beta>0$ and let $\mu_{n}\rightharpoonup\mu$, $\nu_{n}\rightharpoonup\nu$ with respect to weak convergence for $\mu_{n},\nu_{n},\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{d}\times\mathbb{R}^{m})$. Then there exists a subsequence of optimal plans $\alpha_{n_{k}}$ for $W_{2,d_{\beta}}(\mu_{n_{k}},\nu_{n_{k}})$ and an optimal plan $\alpha\in P_{2}\left((\mathbb{R}^{d}\times\mathbb{R}^{m})^{2}\right)$ for $W_{2,d_{\beta}}(\mu,\nu)$ such that $\alpha_{n_{k}}\rightharpoonup\alpha$ weakly.
@@
Let $f:\mathbb{R}^{d}\times\mathbb{R}^{m}\to\mathbb{R}^{d}\times\mathbb{R}^{m}$ be defined by $(y,x)\mapsto(\sqrt{\beta}y,x)$. Then for $\mu_{1},\mu_{2}\in P_{2}(\mathbb{R}^{d}\times\mathbb{R}^{m})$ we have that $W_{2,d_{\beta}}(\mu_{1},\mu_{2})=W_{2}(f_{\sharp}\mu_{1},f_{\sharp}\mu_{2})$ since there is a bijection of couplings
@@
$\int_{(\mathbb{R}^{d}\times\mathbb{R}^{m})^{2}}\|(y_{2},x_{2})-(y_{1},x_{2})\|^{2}\mathrm{d}(f,f)_{\sharp}\alpha=\int_{(\mathbb{R}^{d}\times\mathbb{R}^{m})^{2}}\beta\|y_{2}-y_{1}\|^{2}+\|x_{2}-x_{1}\|\mathrm{d}\alpha$ (108)
@@
which implies that optimal couplings are mapped to optimal couplings. Since also $f_{\sharp}\mu_{n}\rightharpoonup f_{\sharp}\mu,f_{\sharp}\nu_{n}\rightharpoonup f_{\sharp}\nu$ we can use [6, Proposition 7.1.3] to guarantee the existence of a subsequence of optimal plans $\tilde{\alpha}_{n_{k}}$ for $W_{2}(f_{\sharp}\mu_{n_{k}},f_{\sharp}\nu_{n_{k}})$ such that $\tilde{\alpha}_{n_{k}}\rightharpoonup\tilde{\alpha}$ for an optimal plan $\tilde{\alpha}$ for $W_{2}(f_{\sharp}\mu,f_{\sharp}\nu)$. Thus by (107) there exists a subsequence of optimal plans $\alpha_{n_{k}}$ for $W_{2,d_{\beta}}(\mu_{n},\nu_{n})$ such that $\alpha_{n_{k}}\rightharpoonup\alpha$ for an optimal plan $\alpha$ for $W_{2,d_{\beta}}(\mu,\nu)$. ∎
Remark 18.
A useful well known observation is the following. Let $\alpha_{n}\in\Gamma(\mu,\nu),n\in\mathbb{N}$ where $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{l})$. Assume that we have weak convergence $\alpha_{n}\rightharpoonup\alpha$ for some $\alpha\in\mathcal{P}(\mathbb{R}^{2l})$. Then $\alpha\in\Gamma(\mu,\nu)$ and for any measurable function $f:\mathbb{R}^{2l}\to\mathbb{R}$ such that $|f(x_{1},x_{2})|\leq\|x_{1}\|^{2}+\|x_{2}\|^{2}$ we have that $\int_{\mathbb{R}^{2l}}f\mathrm{d}\alpha_{n}\to\int_{\mathbb{R}^{2l}}f\mathrm{d}\alpha$. The latter claim follows from [6, Remark 5.2.3] which implies that $\|x_{1}\|^{2}+\|x_{2}\|^{2}$ is uniformly integrable w.r. to $\{\alpha_{n},n\in\mathbb{N}\}$ since it has fixed marginals $\mu,\nu$ with finite second moments. Then [6, Lemma 5.1.7], which states that $\lim_{n\to\infty}\int f\mathrm{d}\alpha_{n}\to\int f\mathrm{d}\alpha$ for uniformly integrable $f$ and weak converging $\alpha_{n}\rightharpoonup\alpha$, implies the latter claim. It is immediate that $\alpha\in\Gamma(\mu,\nu)$ since $\Gamma(\mu,\nu)$ is compact and thus closed in the weak topology by [6, Remark 5.2.3].
Note that [31, Proposition 3.11] states the following proposition only under some regularity conditions on $\mu,\nu$ which ensures uniqueness of optimal plans. But this is not needed in their proof if one is only interested in the existence of a optimal limit point. For the convenience of the reader we include a proof adapted to our situation but claim no originality.
Proposition 19.
Let $\mu,\nu\in\mathcal{P}_{2,Y}(\mathbb{R}^{d}\times\mathbb{R}^{m})$ and let $\beta_{k}\in\mathbb{R}^{\mathbb{N}}$ be a monotonly increasing series with $\beta_{k}\to\infty$. For a choice of optimal plans $\alpha^{\beta_{k}}$ for $W_{2,d_{\beta_{k}}}(\mu,\nu)$ the the closure of the set $\{\alpha^{\beta_{k}}\}_{k\in\mathbb{N}}$ in $\Gamma(\mu,\nu)$ is compact with respect to the weak convergence topology and every accumulation point $\alpha$ is a optimal plan for $W_{2,Y}(\mu,\nu)$.
Proof.
By [6, Remark 5.2.3] the set $\Gamma(\mu,\nu)$ is compact with respect to the weak convergence topology and thus also the closure of $\{\alpha^{\beta_{k}}:k\in\mathbb{N}\}$ is compact. Consequently accumulation points exist. Let $\alpha\in\Gamma(\mu,\nu)$ be an accumulation point and by abuse of notation let $\alpha^{\beta_{k}}\rightharpoonup\alpha\in\Gamma(\mu,\nu)$. Then by Remark 18 also
$\int_{\mathbb{R}^{2d+2m}}\|y_{1}-y_{2}\|^{2}\mathrm{d}\alpha^{\beta_{k}}\to\int_{\mathbb{R}^{2d+2m}}\|y_{1}-y_{2}\|^{2}\mathrm{d}\alpha=\int_{\mathbb{R}^{2d}}\|y_{1}-y_{2}\|^{2}\mathrm{d}\pi^{1,3}_{\sharp}\alpha.$ (109)
Since the limes on the left is 00 by Proposition 10 we know that $\pi^{1,3}_{\sharp}\alpha$ is only supported on the diagonal. Thus for any $\pi^{1,3}_{\sharp}\alpha$ measurable bounded function $f:\mathbb{R}^{2d}\to\mathbb{R}$ we have that $f=f\circ\Delta\circ\pi^{1}$ for $\pi^{1,3}_{\sharp}\alpha$ a.e. $(y_{1},y_{2})\in\mathbb{R}^{2d}$ and hence
$\int_{\mathbb{R}^{2d}}f\mathrm{d}\pi^{1,3}_{\sharp}\alpha=\int_{\mathbb{R}^{2d}}f\circ\Delta\circ\pi^{1}\mathrm{d}\pi^{1,3}_{\sharp}\alpha=\int_{\mathbb{R}^{2d}}f\mathrm{d}\Delta_{\sharp}P_{Y}$ (110)
which implies $\pi^{1,3}_{\sharp}\alpha=\Delta_{\sharp}P_{Y}$ i.e. $\alpha\in\Gamma^{4}_{Y}(\mu,\nu)$. Furthermore note that $W_{2,\beta}\leq W_{2,Y}$ since every admissible plan for $W_{2,Y}$ is also admissible for $W_{2,d_{\beta}}$ with equal costs and thus
$\int_{\mathbb{R}^{2d+2m}}\|(y_{1},x_{1})-(y_{2},x_{2})\|^{2}\mathrm{d}\alpha^{\beta_{k}}\leq W_{2,Y}(\mu,\nu)^{2}$
for all $k\in\mathbb{N}$. Using Remark 18 we can conclude that
$\int_{\mathbb{R}^{2d+2m}}\|(y_{1},x_{1})-(y_{2},x_{2})\|^{2}\mathrm{d}\alpha=\lim_{k\to\infty}\int_{\mathbb{R}^{2d+2m}}\|(y_{1},x_{1})-(y_{2},x_{2})\|^{2}\mathrm{d}\alpha^{\beta_{k}}\leq W_{2,Y}(\mu,\nu)^{2}.$
Hence $\alpha$ is an optimal plan for $W_{2,Y}(\mu,\nu)$ and thus the claim. ∎
Proof of Proposition 12. By [13, Theorem 8.3.2] we know that weak convergence of probability measures is metrizable and we denote by $d_{weak}$ a metric on $\mathcal{P}\left(\mathbb{R}^{2d}\times\mathbb{R}^{2m}\right)$ that metrizes weak convergence. Let $\alpha_{n}$ be a an optimal plan for $\mu_{n},\nu_{n}$ and $W_{2,d_{\beta_{k}}}(\mu_{n},\nu_{n})$. Then by [6, Proposition 7.1.3] and Lemma 17, there exists a subsequence of $\alpha_{n}$ optimal for $W_{2,d_{\beta_{k}}}(\mu_{n},\nu_{n})$ converging weakly to an optimal plan $\alpha^{\beta_{k}}$ for $W_{2,d_{\beta_{k}}}(\mu,\nu)$. Thus we can find a sequence of optimal plans $\alpha_{n_{k}}$ for $W_{2,d_{\beta_{k}}}(\mu_{n_{k}},\nu_{n_{k}})$ such that $d_{weak}(\alpha^{\beta_{k}},\alpha_{n_{k}})<\frac{1}{k}$. We know by [31, Proposition 3.11] resp. Proposition 19 that $\alpha_{\beta_{k}}\to\alpha$ with respect to $d_{weak}$ for an optimal plan $\alpha\in\Gamma^{4}_{Y}(\mu,\nu)$ for $W_{2,Y}(\mu,\nu)$. Thus, for $\epsilon>0$, there exists a $k$ such that $\frac{1}{k}+d_{weak}(\alpha^{\beta_{k}},\alpha)<\epsilon$ and we obtain
$d_{weak}(\alpha_{n_{k}},\alpha)\leq d_{weak}(\alpha_{n_{k}},\alpha^{\beta_{k}})+d_{weak}(\alpha^{\beta_{k}},\alpha)\leq\frac{1}{k}+d_{weak}(\alpha^{\beta_{k}},\alpha)<\epsilon$
@@
Appendix F Implementation Details
We use a setup similar to [55], using the time dependent U-Net architecture from [46] which are trained using Adam [36]. As in [55] we clip the gradient norm to 1 and use exponential moving averaging with a decay of $0.9999$. The differences are we use a constant learning rate of 2e-4, 256 model channels and no dropout. We train using 50k target samples for 500 epochs using a batch size of 500 for the minibatch OT couplings and a batch size of 100 for training the networks. We set the same random seed during training to be able to compare runs for different sources of couplings. The conditional coupling plans are calculated using the Python Optimal Transport package [22]. For inference simulate the corresponding ODEs using the torchiffeq [17] package. To evaluate our results, we use the Fréchet inception distance (FID) [30] [^2] We use the implementation from https://github.com/mseitzer/pytorch-fid.. We compute the distance on 50k training samples, for which we generate 50k samples given the same labels as the training samples.